# latexFont shape

### v1

@@ Abstract
In this paper, we consider the problem of referring camouflaged object detection (Ref-COD), a new task that aims to segment specified camouflaged objects based on some form of reference, e.g., image, text. We first assemble a large-scale dataset, called R2C7K, which consists of 7K images covering 64 object categories in real-world scenarios. Then, we develop a simple but strong dual-branch framework, dubbed R2CNet, with a reference branch learning common representations from the referring information and a segmentation branch identifying and segmenting camouflaged objects under the guidance of the common representations. In particular, we design a Referring Mask Generation module to generate pixel-level prior mask and a Referring Feature Enrichment module to enhance the capability of identifying camouflaged objects. Extensive experiments show the superiority of our Ref-COD methods over their COD counterparts in segmenting specified camouflaged objects and identifying the main body of target objects. Our code and dataset are publicly available at https://github.com/zhangxuying1004/RefCOD.
@@
Referring Camouflaged Object Detection; Referring Images; R2C7K Dataset; R2CNet Framework.
@@ 1 Introduction
Camouflaged object detection (COD), which aims to segment objects that are visually hidden in their surroundings, has been attracting more and more attention [1, 2, 3, 4, 2]. This research topic plays an important role in a wide range of real-world applications, e.g., medical image segmentation [5, 6], surface defect detection [7], and pest detection [8]. It is noteworthy that although there may be multiple camouflaged objects in a real scene, a large number of applications hope to find the specified ones. A typical example should be the explorers are looking for some special species but most of them may be hidden deep together with other similar objects. In this case, if we have some references about the targets, the finding process will become orientable and thus get easier. As a result, it is promising to explore the COD with references, which is abbreviated as Ref-COD in this paper.
CODRef-COD(Ours)ImageCOD-GTRefCOD-GT
@@
Figure 1: Visual comparison between the standard COD and our Ref-COD. Given an image containing multiple camouflaged objects, the COD model tends to detect all possible camouflaged objects that are blended into the background without discrimination, while the Ref-COD model attempts to identify the camouflaged objects under the condition of some form of reference. Note that ‘Ref1’ and ‘Ref2’ are two forms of reference, where the former denotes a text description about the target objects and the latter denotes an image with salient target objects.
@@
Fig. 1 illustrates the task relationship between the standard COD and our Ref-COD. In particular, our Ref-COD leverages the referring information to guide the identification of specified camouflaged objects, which is consistent with that of human visual perception to camouflaged objects [9]. It transforms COD from searching aimlessly for differential regions in camouflage scenes to matching target objects with specified purposes, and the key issue lies on which form of information is appropriate to be used as a reference. Recent works have explored several forms of referring information for image segmentation, e.g., referring expression segmentation with text reference [10], and few-shot segmentation with image reference [11]. However, whether the annotated images containing specified camouflaged objects or detailed textual descriptions of camouflaged objects for existing images, the acquisition process is time-consuming and laborious, which hinders the transfer of these methods to COD. Considering that images with salient objects are readily available on the Internet, the common representations of specified objects can also be obtained from them with advanced research on foreground prediction. Besides, the increasingly popular CLIP and prompt engineering [12] also make simple text expressions become powerful referring information for specified objects.
@@
Based on the aforementioned two references, we propose a novel Ref-COD benchmark. To enable a comprehensive study on this new benchmark, we build a large-scale dataset, named R2C7K, which contains a large number of samples without copyright disputes in real-world scenarios. And the outline of this dataset is as follows: 1) It has 7K images covering 64 object categories; 2) It consists of two subsets, i.e., the Camo-subset composed of images containing camouflaged objects and the Ref-subset composed of images containing salient objects; 3) The number of images for each category in the Ref-subset is fixed, while the one in the Camo-subset is not. 4) The expression ‘a photo of [CLASS]’ forms the textual reference for specified objects.
@@
To investigate the role of the referring information in Ref-COD, we design a dual-branch network architecture and develop a simple but effective framework, named R2CNet. This framework includes a reference branch and a segmentation branch. The reference branch aims to capture common feature representations of specified objects from the referring images composed of salient objects or textual descriptions, which will be used to retrieve the target objects. Particularly, we build a Referring Mask Generation (RMG) module to generate pixel-level referring information. In this module, a dense comparison is performed between the common representations from the reference branch and each position of the visual features from the segmentation branch to generate a referring prior mask. However, there may exist variances in appearance between the camouflaged objects and the salient objects even though they belong to the same category, as well as modal differences between text expressions and images, which may increase the difficulty of retrieving accurate camouflaged objects. To overcome this shortcoming, a dual-source information fusion motivated by multi-modal fusion is employed to eliminate the information differences between two information sources. In addition, we also design a Referring Feature Enrichment (RFE) module to achieve the interaction among multi-scale visual features, and further highlight the target objects.
@@
  • We propose a new benchmark, termed Ref-COD, which, to the best of our knowledge, is the first attempt to directionally segment the camouflaged objects with simple referring information.
@@ 2 Related Work
In this section, we first present the research progress and existing problems of COD tasks. Then, we introduce the development history of the SOD topic. Finally, we describe referring object segmentation methods with different forms of reference information.
@@ 2.1 Camouflaged Object Detection
Identifying camouflaged objects from a given scene is challenging due to the high similarity to their surroundings, diversity in scale, fuzziness in appearance, etc. To solve this task, an increasing number of works have been proposed. One of the early systematic studies is presented by [18], which publishes a large-scale dataset with high-quality annotation, namely COD10K, and a well-designed search and identify framework for this research. Later on, a large number of strategies, e.g., multi-scale feature fusion [19, 20, 21, 2, 22], multi-stage refinement [23, 1, 24, 25], graph learning [26], weakly supervised [27], uncertainty [28, 29, 30], foreground and background separation [31, 32, 33] , and attention mechanism [34, 35] have been proposed to improve the performance of the COD models. More related works are presented in recent survey paper [36]. More recently, numerous works introduce additional information, e.g., boundary [37, 4, 38, 39, 40], texture [41, 42, 3], frequency [43, 44] and , depth [45, 46, 47, 47], as guidance to further improve the accuracy of the segmentation results. However, obtaining these forms of information in high quality from the existing camouflaged images requires expensive time and resource costs, and they cannot explicitly tell the COD model what to segment.
In this paper, we propose a novel benchmark, which extracts common representations of a certain category from readily available referring images as explicit semantic guidance to locate and segment the camouflaged objects.
@@ 2.2 Salient Object Detection
Salient object detection aims to capture the most attention-grabbing objects in the given image. And the development of this topic has witnessed two stages: 1) the traditional method stage and 2) the deep learning method stage. During the former period, hand-crafted features [48, 49, 50, 51] from image patches, object proposals, and super-pixels, play the most important role. Although these methods work well in some scenes, the extraction of these features is time-consuming and their performance may sharply degrade in complex conditions. With the prosperity of fully convolutional networks [52] and Transformers [53, 54, 55, 56], SOD research enters the latter stage. In this stage, U-shape structure [13, 57], multi-stage supervision [58, 59, 60], and attention mechanisms [61, 62] are widely in SOD methods to achieve pixel-wise predictions.
It is worth mentioning that SOD has extensive applications in various research communities, e.g., visual tracking [63], unsupervised segmentation [64], image compression [65], and content-aware image editing [66], mainly because it helps to find the objects or regions that efficiently represent a scene. And this property of SOD also inspires the reference idea of this paper.
Ref-SubsetCamo-Subset
Cat
Butterfly
Bird
@@ 2.3 Referring Object Segmentation
Referring Object Segmentation means segmenting visual objects from a given image under a certain form of references, e.g., image, text, etc.
Few-shot segmentation (FSS) explores object segmentation guided by annotated images containing objects of the same category, where the model is trained on a large number of images whose pixels are labeled with base classes (query set) and performs dense pixel prediction on unseen classes given a few annotated samples (support set). Particularly, most existing FSS networks include two branches, i.e., support branch and query branch, to extract the features of support images and query images and achieve the interaction between them. The pioneering work of FSS research is proposed by [67], where the support branch directly predicts the weights of the last layer in the query branch for segmentation. Then, the masked average pooling operation is proposed by [68] to extract representative support features, which is widely adopted by subsequent works. More recently, a large number of works [11, 69, 69] build powerful modules on the frozen backbone network to improve the adaptability of the models to unseen categories.
Referring Expression Segmentation (RES) explores object segmentation guided by a text expression. RES aims to segment visual objects based on a given textual expression, and the two-branch architecture is also adopted by the networks in this research. The first work is introduced by [10], where the visual and linguistic features are first extracted by a visual encoder and a language encoder respectively, and their concatenation is employed to generate the segmentation mask. Subsequently, a series of methods based on multi-level visual features [70], multi-modal LSTM [71], attention mechanism [72, 73], collaborative network [74] are incorporated in RES methods successively to generate more accurate results. In addition, the text descriptions are also adopted by [75] as references for the richness of image content to achieve a better fixation prediction.
With the rise of CLIP and prompt engineering, a series of works [76, 77] adopt image or text prompts to facilitate vision tasks. It is noteworthy that such prompts can also be regarded as reference information. For example, CLIPSeg [76] employs visual or text prompts from the pre-trained CLIP model as references to address image segmentation.
In this paper, the proposed Ref-COD also belongs to a referring object segmentation task. However, different from the existing methods, the collection of its referring information does not take much effort. To be specific, it neither needs to collect rare and hard-to-label images containing camouflaged objects of the same category nor annotates detailed text descriptions for existing COD datasets, which is convenient for academia and industry to follow.
@@ 3.2 Data Statistics
Subset Comparisons. Fig. 3 presents 4 attribute comparisons between the images in Ref-subset and Camo-subset. Specifically, the object area refers to the size of the objects in a given image, the object ratio is the proportion of objects in an image, the object distance is the distance from the object center to the image center, and the global contrast is a metric to evaluate how challenging an object is to detect. It can be observed that the objects in the Ref-subset are bigger than those objects in the Camo-subset, and the images in the Ref-subset contain more contrast cues. Therefore, the objects in the Ref-subset are easier to be detected, which means that this form of referring information is readily available and suitable for the Ref-COD research.
@@
Resolution Distribution. Fig. 5(a) and Fig. 5(b) shows the resolution distribution of the images in the Camo-subset and Ref-subset. As can be seen, these two subsets contain a large number of Full HD images, which can provide more details on the object boundaries and textures.
Dataset Splits. To facilitate the development of models for Ref-COD research, we provide a referring split for the R2C7K dataset. For the Ref-subset, 20 samples are randomly selected from each category for training while the remaining samples in each category are used for testing; As far as the Camo-subset, the samples coming from the training set of COD10K are also utilized for training and the ones belonging to the test set are used for testing. And the samples from NC4K are randomly assigned to the training and testing sets to ensure that each category in these two splits contains at least 6 samples.
Figure 6: Overall architecture of our R2CNet framework, which is composed of two branches, i.e., reference branch in green and segmentation branch in orange. In the reference branch, the common representation of a specified object from the textural expression is extracted by the text encoder of CLIP, while the one from images is obtained by masking and pooling the visual features with the foreground map generated by the foreground prediction network. In the segmentation branch, the visual features from the last three layers of the encoder are employed to represent the given image. Then, these two kinds of feature representations are fused and compared in the well-designed RMG module to generate a mask prior, which is used to enrich the visual feature at different scales to highlight the camouflaged targets in our RFE module. Finally, the enriched features are fed into the decoder to generate the final segmentation map.
@@ 4 Proposed Framework
In this section, we first briefly describe the definition of our Ref-COD task in Sec. 4.1. Then, we introduce the overall architecture of our R2CNet in Sec. 4.2. Next, we present the details of the Referring Mask Generation (RMG) module and the Referring Feature Enrichment (RFE) module in Sec. 4.3 and Sec. 4.4 respectively.
@@ 4.1 Task Description
Ref-COD aims at segmenting the camouflaged objects under the guidance of a textual description or a few referring images. In particular, the input of a Ref-COD system is composed of two parts, where the first part is an image with camouflaged objects, termed as $I^{camo}\in\mathbb{R}^{H\times W\times 3}$, while the other part can be either a textual expression, denoted as $E^{ref}=\{e_{i}\}_{i=1}^{4}$, or several referring images with salient objects of a certain category, denoted as $I^{ref}=\{I^{ref}_{i}\}_{i=1}^{K},I^{ref}_{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the height and width of given images, $e_{i}$ is a word in the textual expression, and $K$ represents the number of referring images. Note that $I^{camo}$ comes from the Camo-subset of R2C7K, and it contains camouflaged objects in a specified category c. Meanwhile, $I^{ref}_{i}$ is sampled from the Ref-subset, and its salient objects are in category c. And the output of Ref-COD is the binary mask $M^{seg}$ for the camouflaged objects of category c in $I^{camo}$.

### v3

@@ Abstract
We consider the problem of referring camouflaged object detection (Ref-COD), a new task that aims to segment specified camouflaged objects based on a small set of referring images with salient target objects. We first assemble a large-scale dataset, called R2C7K, which consists of 7K images covering 64 object categories in real-world scenarios. Then, we develop a simple but strong dual-branch framework, dubbed R2CNet, with a reference branch embedding the common representations of target objects from referring images and a segmentation branch identifying and segmenting camouflaged objects under the guidance of the common representations. In particular, we design a Referring Mask Generation module to generate pixel-level prior mask and a Referring Feature Enrichment module to enhance the capability of identifying specified camouflaged objects. Extensive experiments show the superiority of our Ref-COD methods over their COD counterparts in segmenting specified camouflaged objects and identifying the main body of target objects. Our code and dataset are publicly available at https://github.com/zhangxuying1004/RefCOD.
@@
Referring Camouflaged Object Detection; Common Representations; R2C7K Dataset; R2CNet Framework.
@@ 1 Introduction
Camouflaged object detection (COD), which aims to segment objects that are visually hidden in their surroundings, has been attracting more and more attentions [1, 2, 3, 4]. This research topic plays an important role in a wide range of real-world applications, e.g., medical image segmentation [5, 6], surface defect detection [7, 8], and pest detection [9]. It is noteworthy that although there may be multiple camouflaged objects in a real scene, we just want to find the specified ones in many applications. A typical example should be the explorers are looking for some special species but most of them may be hidden deep together with other similar objects. In this case, if we have some references about the targets, the finding process will become orientable and thus get easier. As a result, it is promising to explore the COD research with references, which is abbreviated as Ref-COD in this paper.
[…]
(a) Standard COD
[…]
(b) Ref-COD (Ours)
@@
Figure 1: Visual comparison between the standard COD and our Ref-COD. Given an image containing multiple camouflaged objects, the COD model tends to find all possible camouflaged objects that are blended into the background without discrimination, while the Ref-COD model attempts to identify the specified camouflaged objects under the condition of a set of referring images.
@@
Ref-COD leverages the referring information to guide the identification of specified camouflaged objects, which is consistent with that of human visual perception to camouflaged objects [10]. And the key issue lies on which form of information is appropriate to be used as a reference. Recent works have explored several forms of reference for image segmentation, e.g., referring expression segmentation with text reference [11], and few-shot segmentation with image reference [12]. However, whether the annotated images containing specified camouflaged objects or detailed textual descriptions of camouflaged objects for existing images, the acquisition process is time-consuming and laborious, which hinders the transfer of these methods to COD. Considering that images with salient objects are readily available on the Internet, a straightforward question arises: Is it possible for us to take advantage of such images to help better identify specified camouflaged objects?
@@
Motivated by this question, we propose a novel Ref-COD benchmark. Our intention is to leverage the increasingly advanced research on salient object detection (SOD) to acquire common representations of target objects from referring images, which are used to guide the segmentation of specified camouflaged objects. Fig. 1 illustrates the task relationship between the standard COD and our Ref-COD. In particular, our Ref-COD transforms the process of COD from detecting aimlessly the differential objects in camouflage scenes to matching target objects with specified purposes. To enable a comprehensive study on this new benchmark, we build a large-scale dataset, named R2C7K, which contains a large number of samples without copyright disputes in real-world scenarios. The basic information of this dataset is as follows: 1) It has 7K images covering 64 object categories; 2) It consists of two subsets, i.e., the Camo-subset composed of images containing camouflaged objects and the Ref-subset composed of images containing salient objects; 3) The number of images for each category in the Ref-subset is fixed, while the one in the Camo-subset is not.
@@
To investigate the role of the referring information in Ref-COD, we design a dual-branch network architecture and develop a simple but effective framework, named R2CNet. This framework includes a reference branch and a segmentation branch. The reference branch aims to capture common representations of target objects from the referring images composed of salient objects, which will be used to identify the specified camouflaged objects. Particularly, we build a Referring Mask Generation (RMG) module to generate pixel-level referring information. In this module, a dense comparison is performed between the common representations from the reference branch and each position of the visual features from the segmentation branch to generate a referring prior mask. However, there may exist variances in appearance between the camouflaged objects and the salient objects even though they belong to the same category, which may increase the difficulty of identifying accurate camouflaged objects. To overcome this shortcoming, a dual-source information fusion strategy is employed to eliminate the information differences between two information sources. In addition, we also design a Referring Feature Enrichment (RFE) module to achieve the interaction among multi-scale visual features under the guidance of referring mask, and further highlight the target objects.
@@
  • We propose a new benchmark, termed Ref-COD. To the best of our knowledge, it is the first attempt to bridge SOD and COD and segment the specified camouflaged objects with salient object images.
@@ 2 Related Work
In this section, we first present the research progress and existing problems of COD tasks. Then, we introduce the development history of the SOD topic. Finally, we describe referring object segmentation research with different forms of reference information.
@@ 2.1 Camouflaged Object Detection
Identifying camouflaged objects from a confusing scene is challenging due to the high similarity to their surroundings, diversity in scale, fuzziness in appearance, etc. To solve this task, an increasing number of works have been proposed. One of the early systematic studies is presented by [18], which publishes a large-scale dataset with high-quality annotation, namely COD10K, and a well-designed search and identify framework for this research. Later on, a large number of strategies, e.g., multi-scale feature fusion [19, 20, 21, 2, 22], multi-stage refinement [23, 1, 24, 25], graph learning [26], weakly supervised [27], uncertainty [28, 29, 30], foreground and background separation [31, 32, 33] , and attention mechanism [34, 35] have been proposed to improve the performance of the COD models. More related works are presented in recent survey paper [36]. More recently, numerous works introduce additional information, e.g., boundary [37, 4, 38, 39, 40], texture [41, 42, 3], frequency [43, 44], and depth [45, 46, 47, 47], as guidance to further improve the accuracy of the segmentation results. However, obtaining these forms of information of high quality from the existing camouflaged images requires expensive time and resource costs, and they cannot explicitly tell the COD model what to segment.
In this paper, we propose a novel benchmark, which extracts common representations of target objects from readily available referring images as explicit semantic guidance to locate and segment the specified camouflaged objects. This make our work quite different from the standard COD task.
[…]
@@ 2.2 Salient Object Detection
Salient object detection (SOD) aims to capture the most attention-grabbing objects in the given image. The development of this topic has witnessed two stages: 1) the traditional method stage and 2) the deep learning method stage. During the former period, hand-crafted features [48, 49, 50, 51] from image patches, object proposals, and super-pixels, play the most important role. Although these methods work well in some scenes, the extraction of these features is usually time-consuming. More importantly, their performance may sharply degrade in complex conditions. With the prosperity of fully convolutional networks [52] and Transformers [53, 54, 55, 56, 57], SOD research enters the latter stage. In this stage, U-shape structure [13, 58], multi-stage supervision [59, 60, 61], and attention mechanisms [62, 63] are widely used in SOD methods to achieve more accurate pixel-wise predictions.
It is worth mentioning that SOD has extensive applications in various research communities, e.g., visual tracking [64], unsupervised segmentation [65], image compression [66], and content-aware image editing [67], mainly because it can help find the objects or regions that can efficiently represent a scene. This property of SOD also inspires the reference idea of this paper.
@@ 2.3 Referring Object Segmentation
Referring Object Segmentation means segmenting visual objects from a given image under a certain form of references, e.g., image and text.
Few-shot segmentation (FSS) explores object segmentation guided by annotated images containing objects of the same category. The models are trained on a large number of images whose pixels are labeled with base classes (query set) and perform dense pixel prediction on unseen classes given a few annotated samples (support set). Particularly, most existing FSS networks include two branches, i.e., support branch and query branch, to extract the features of support images and query images and achieve the interaction between them. The pioneering work of FSS research is proposed by [68], where the support branch directly predicts the weights of the last layer in the query branch for segmentation. Then, the masked average pooling operation is proposed by [69] to extract representative support features, which is widely adopted by subsequent works. More recently, a large number of works [12, 70, 71] build powerful modules on the frozen backbone network to improve the adaptability of the models to unseen categories.
Referring Expression Segmentation (RES) explores object segmentation guided by a textual expression. RES aims to segment visual objects based on the expression, and the two-branch architecture is also adopted by the networks in this research. The first work is introduced by [11], where the visual and linguistic features are first extracted by a visual encoder and a language encoder respectively, and their concatenation is employed to generate the segmentation mask. Subsequently, a series of methods based on multi-level visual features [72], multi-modal LSTM [73], attention mechanism [74, 75], collaborative network [76] are incorporated in RES methods successively to generate more accurate results. In addition, the text descriptions are also adopted by [77] as references for the richness of image content to achieve a better fixation prediction.
In this paper, the proposed Ref-COD also belongs to a referring object segmentation task. However, different from the existing methods, the collection of its referring information does not take much effort. To be specific, it neither needs to collect rare and hard-to-label images containing target camouflaged objects nor annotates detailed text descriptions for existing COD datasets, which is convenient for academia and industry to follow.
@@ 3.2 Data Statistics
Subset Comparisons. Fig. 3 presents four attribute comparisons between the images in Ref-subset and Camo-subset. Specifically, the object area refers to the size of the objects in a given image, the object ratio is the proportion of objects in an image, the object distance is the distance from the object center to the image center, and the global contrast is a metric to evaluate how challenging an object is to detect. It can be observed that the objects in the Ref-subset are bigger than those objects in the Camo-subset, and the images in the Ref-subset contain more contrast cues. Therefore, the objects in the Ref-subset are easier to be detected, and this form of referring information is suitable for the Ref-COD research.
@@
Figure 6: Overall architecture of our R2CNet framework, which is composed of two branches, i.e., reference branch in green and segmentation branch in orange. In the reference branch, the common representations of a specified object from images is obtained by masking and pooling the visual features with the foreground map generated by a SOD network. In the segmentation branch, the visual features from the last three layers of the encoder are employed to represent the given image. Then, these two kinds of feature representations are fused and compared in the well-designed RMG module to generate a mask prior, which is used to enrich the visual features among different scales to highlight the camouflaged targets in our RFE module. Finally, the enriched features are fed into the decoder to generate the final segmentation map.
TABLE I: Comparisons of our R2C7K with previous COD datasets across a wide range of dimensions. Cate.: the number of categories. Camo-Img.: the number of images containing camouflaged objects; Ref-Img.: the number of referring images; Loc.: Location; Det.: Detection; Cls.: Classification; WS.: Weak Supervision; RefSeg.: Referring Object Segmentation.
Datasets Statistics Tasks
Year Cate. Camo-Img. Ref-Img. Loc. Det. Cls. WS. RefSeg.
CHAMELEON [83] 2018 N/A 76 N/A
CAMO [80] 2019 N/A 2500 N/A
COD10K [1] 2020 68 5066 N/A
NC4K [81] 2021 N/A 4121 N/A
R2C7K (Ours) 2023 64 5015 1600
@@
Resolution Distribution. Fig. 5a and Fig. 5b shows the resolution distribution of the images in the Camo-subset and Ref-subset. As can be seen, these two subsets contain a large number of Full HD images, which can provide more details on the boundaries and textures of objects.
Dataset Splits. To facilitate the development of models for Ref-COD research, we provide a referring split for the R2C7K dataset. For the Ref-subset, 20 samples are randomly selected from each category for training while the remaining 5 samples in each category are used for testing; As for the Camo-subset, the samples coming from the training set of COD10K are also utilized for training and the ones belonging to the test set are used for testing. The samples from NC4K are randomly assigned to the training and testing sets to ensure that each category in these two splits contains at least 6 samples.
Comparison with Existing Datasets. We compare the proposed R2C7K dataset with the existing COD datasets, including CHAMELEON [83], CAMO [80], COD10K [1], and NC4K [81], across a wide range of dimensions. As shown in Tab. I, the CHAMELEON dataset, collected from Google search engine with a keyword, i.e., “concealed animal”, only contains 76 images. The CAMO dataset has 2,500 images containing camouflaged objects across 8 categories. The COD10K and NC4K datasets have a higher number of images containing camouflaged objects, with 5,066 and 4,121 images respectively. These datasets greatly promote the development of camouflage object detection. Nevertheless, they still have limitations when it comes to a real scene with multiple types of camouflaged objects, hindering their practical application. In contrast, our R2C7K dataset has a large number of simple scene images (1,600) covering multiple object categories (64) so that users can choose them as references according to their needs to search for specified camouflaged objects in complex scenes (5,015). As a result, it has the potential to benefit more tasks.
@@ 4 Proposed Framework
In this section, we first briefly describe the definition of our Ref-COD task in Sec. 4.1. Then, we introduce the overall architecture of our R2CNet in Sec. 4.2. Next, we present the details of the Referring Mask Generation (RMG) module and the Referring Feature Enrichment (RFE) module in Sec. 4.3 and Sec. 4.4, respectively.
@@ 4.1 Task Description
Our Ref-COD aims at segmenting the specified camouflaged objects under the guidance of several referring images. Different from the standard COD that detects all the camouflage objects, our Ref-COD attempts to identify the specific objects by matching the targets provided by the references. In particular, the input of a Ref-COD system is composed of two parts, where the first part is an image with camouflaged objects, termed as $I^{camo}\in\mathbb{R}^{3\times H\times W}$, while the other part is a few referring images with salient target objects, denoted as $I^{ref}=\{I^{ref}_{i}\}_{i=1}^{K},I^{ref}_{i}\in\mathbb{R}^{3\times H\times W}$, where $H$ and $W$ denote the height and width of the given images, and $K$ represents the number of referring images. Note that $I^{camo}$ comes from the Camo-subset of R2C7K, and it contains camouflaged objects in a specified category c. Meanwhile, $I^{ref}_{i}$ is sampled from the Ref-subset, and its salient objects are in category c. The output of Ref-COD is a binary mask $M^{seg}$ for the camouflaged objects of category c in $I^{camo}$.